two cards are drawn from an ordinary deck of 52 cards. Find the probability of
a. two red aces
b. two honor cards ( A K Q J 10)

Answer by Alan3354(69443)

two cards are drawn from an ordinary deck of 52 cards. Find the probability of
a. two red aces
Assuming no replacement:
The 1st draw has a chance of 2 in 52 = 1 in 26.
The 2nd draw has a chance of 1 in 51.
Overall, it's 1 in 26*51 = 1 in 1326.
----------------------------
b. two honor cards ( A K Q J 10)
The 1st draw has a chance of 20/52 = 5/13.
The 2nd draw has a chance of 19/51.
20*19/13*51 = 380/663